find 2 positive numbers whoes squares have a sum of 113 and a difference of 15
Answer by stanbon(57307)

find 2 positive numbers whoes squares have a sum of 113 and a difference of 15
-----------
EQUATIONS:
x^2 + y^2 = 113
x^2 - y^2 = 15
----------------
Subtract the 2nd from the 1st to get:
2y^2 = 98
y^2 = 49
y = 7
---------
Substitute to solve for "x":
x^2 + 49 = 113
x^2 = 64
x = 8
===========
Cheers,
Stan H.